# Nash-equilibrium Seeking Algorithm for Power-Allocation Games on Networks of International Relations

### v1

@@ 3.2 Convergence analysis
If $u_{i_{t_{0}}}(\hat{X}^{(3)}_{0})\geq u_{i_{t_{0}}}(X^{(3)}_{0})+1$, there exists an available power-allocation strategy for country $i_{t_{0}}$ to strictly increase its utility when it is selected to update at time $1$, which contradicts the property of the set $\mathbf{S_{2}}$. Hence $u_{i_{t_{0}}}(\hat{X}^{(3)}_{0})\leq u_{i_{t_{0}}}(X^{(3)}_{0})$. Let $u^{\triangle}=u_{i_{t_{0}}}(X^{(3)}_{0})-u_{i_{t_{0}}}(\hat{X}^{(3)}_{0})$.
@@
So the reason that country $i_{t_{0}}$ can not increase its utility by applying the status-quo-breaking strategy at time 0 is that at least $u^{\triangle}+1$ country $i_{t_{0}}$’s friendly countries transfers from safety to danger if country $i_{t_{0}}$ takes the status-quo-breaking strategy at time 0, but not being dangerous given the strategy matrix $X^{(3)}_{t_{0}}$. Denote by $i_{f_{1}},i_{f_{2}},...,i_{f_{u^{\triangle}+1}}$ this kind of country $i_{t_{0}}$’s friendly countries. Let $t^{*}_{k}:=x^{(3)}_{i_{t_{0}}i_{f_{k}}}(0)-x^{(3)}_{i_{t_{0}}i_{f_{k}}}({t_{0}})-s_{i_{f_{k}}}(X^{(3)}_{0})$ and $t^{*}=\sum_{k=1}^{u^{\triangle}+1}t^{*}_{k}+1$. We can construct a sequence of strategy matrices $\{\bar{X}^{(3)}_{t}\}_{0\leq t\leq t^{*}}=\mathbf{X}(X^{(3)},\{i_{t}\}_{1\leq t\leq t^{*}})$ which we denote by $\bar{\mathbf{X}}^{(3)}$. Let $i_{t}=i_{f_{k}}$ for any $\sum_{j=1}^{k-1}t^{*}_{j}\leq t\leq\sum_{j=1}^{k}t^{*}_{j}-1$ and all $1\leq k\leq u^{\triangle}+1$. Let $i_{t^{*}}=i_{t_{0}}$. Meanwhile, before time $t^{*}$, let the updating country’s power increase by $1$ and make the updating country to adopt a feasible strategy in which the allocation of power to the other countries remains unchanged, and any additional power is solely used to protect itself. Let $i_{t_{0}}$ choose the status-quo-breaking strategy at time $t^{*}$. Then for all $1\leq k\leq u^{\triangle}+1$, country $i_{f_{k}}$’s state will transfer from safe to precarious during the update at time $t^{*}$. Since the updating countries befor $t^{*}$ are all in the safe state, according to what was explained in step 2, no country’s state will change before $t^{*}$ with the sequence of strategy matrices $\{\bar{X}^{(3)}_{t}\}_{0\leq t\leq t^{*}}$, which means that $\{\bar{X}^{(3)}_{t}\}_{0\leq t\leq t^{*}-1}$ is a feasible sequence. Notice that

### v2

@@ 3.2 Convergence analysis
If $u_{i_{t_{0}}}(\hat{X}^{(3)}_{0})\geq u_{i_{t_{0}}}(X^{(3)}_{0})+1$, there exists an available power-allocation strategy for country $i_{t_{0}}$ to strictly increase its utility when it is selected to update at time 11, which contradicts the property of the set $\mathbf{S_{2}}$. Hence $u_{i_{t_{0}}}(\hat{X}^{(3)}_{0})\leq u_{i_{t_{0}}}(X^{(3)}_{0})$. Let $u^{\triangle}=u_{i_{t_{0}}}(X^{(3)}_{0})-u_{i_{t_{0}}}(\hat{X}^{(3)}_{0})$.
@@
So the reason that country $i_{t_{0}}$ can not increase its utility by applying the status-quo-breaking strategy at time 0 is that at least $u^{\triangle}+1$ country $i_{t_{0}}$’s friendly countries transfers from safety to danger if country $i_{t_{0}}$ takes the status-quo-breaking strategy at time 0, but not being dangerous given the strategy matrix $X^{(3)}_{t_{0}}$. Denote by $i_{f_{1}},i_{f_{2}},...,i_{f_{u^{\triangle}+1}}$ this kind of country $i_{t_{0}}$’s friendly countries. Let $t^{*}_{k}:=x^{(3)}_{i_{t_{0}}i_{f_{k}}}(0)-x^{(3)}_{i_{t_{0}}i_{f_{k}}}({t_{0}})-s_{i_{f_{k}}}(X^{(3)}_{0})$ and $t^{*}=\sum_{k=1}^{u^{\triangle}+1}t^{*}_{k}+1$. We can construct a sequence of strategy matrices $\{\bar{X}^{(3)}_{t}\}_{0\leq t\leq t^{*}}=\mathbf{X}(X^{(3)},\{i_{t}\}_{1\leq t\leq t^{*}})$ which we denote by $\bar{\mathbf{X}}^{(3)}$. Let $i_{t}=i_{f_{k}}$ for any $\sum_{j=1}^{k-1}t^{*}_{j}\leq t\leq\sum_{j=1}^{k}t^{*}_{j}-1$ and all $1\leq k\leq u^{\triangle}+1$. Let $i_{t^{*}}=i_{t_{0}}$. Meanwhile, before time $t^{*}$, let the updating country’s power increase by 11 and make the updating country to adopt a feasible strategy in which the allocation of power to the other countries remains unchanged, and any additional power is solely used to protect itself. Let $i_{t_{0}}$ choose the status-quo-breaking strategy at time $t^{*}$. Then for all $1\leq k\leq u^{\triangle}+1$, country $i_{f_{k}}$’s state will transfer from safe to precarious during the update at time $t^{*}$. Since the updating countries befor $t^{*}$ are all in the safe state, according to what was explained in step 2, no country’s state will change before $t^{*}$ with the sequence of strategy matrices $\{\bar{X}^{(3)}_{t}\}_{0\leq t\leq t^{*}}$, which means that $\{\bar{X}^{(3)}_{t}\}_{0\leq t\leq t^{*}-1}$ is a feasible sequence. Notice that